A boy was offered a bonus if he sold 100 subscriptions to a magazine. Each day he sold three subscriptions more than he had sold on the previous day, and on the eighth day he reached a total of exactly 100 subscriptions. How many subscriptions did he sell each day?
Answer by richwmiller(17219)

y=mx+b
100 8
97 7
94 6
91 5
88 4
85 3
82 2
79 1
y-y=m*(x-x)
82-79=m(2-1)
3=m
79=3*1+b
79=3+b
76=b
y=3x+76
check
100=3*8+76
100=24+76
check